# Optimal Neural Compressors for the Rate-Distortion-Perception Tradeoff

### v1

@@ Abstract
Recent efforts in neural compression have focused on the rate-distortion-perception (RDP) tradeoff, where the perception constraint ensures the source and reconstruction distributions are close in terms of a statistical divergence. Theoretical work on RDP describes interesting properties of RDP-optimal compressors without providing constructive and low complexity solutions. While classical rate distortion theory shows that optimal compressors should efficiently pack the space, RDP theory additionally shows that infinite randomness shared between the encoder and decoder may be necessary for RDP optimality. In this paper, we propose neural compressors that are low complexity and benefit from high packing efficiency through lattice coding and shared randomness through shared dithering over the lattice cells. For two important settings, namely infinite shared and zero shared randomness, we analyze the rate, distortion, and perception achieved by our proposed neural compressors and further show optimality in the presence of infinite shared randomness. Experimentally, we investigate the roles these two components of our design, lattice coding and randomness, play in the performance of neural compressors on synthetic and real-world data. We observe that performance improves with more shared randomness and better lattice packing.
@@ 1 Introduction
$\mathrm{s.t.}$ $\operatorname{\mathbb{E}}_{P_{X}P_{\hat{X}|X}}[\operatorname{\Delta}(X,\hat{X})]\leq D$
$\delta(P_{X},P_{\hat{X}})\leq P,$
@@
(a) Lattice transform coding (LTC); fully deterministic.
(b) Private-dither LTC (PD-LTC); randomness not shared.
(c) Shared-dither LTC (SD-LTC); infinite shared randomness.
(d) Quantized-shared-dither LTC (QSD-LTC) with nested lattices $\Lambda,\Lambda_{f}$; $\frac{1}{n}\log|\Lambda/\Lambda_{f}|$ bits of shared randomness.
@@
Figure 1: Lattice transform coding (LTC) with different amounts of (shared) randomness using dithering; $\bm{u}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda))$ and $\bm{u}_{f}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda_{f}))$ are continuous, and $\hat{\bm{u}}\sim\mathrm{Unif}(|\Lambda/\Lambda_{f}|)$ is discrete. LTC (1(a)) and PD-LTC (1(b)) entropy-code the quantized latent $\hat{\bm{y}}=Q_{\Lambda}(\bm{y})$ with entropy model $p_{\hat{\bm{y}}}$. SD-LTC (1(c)) and QSD-LTC (1(d)) entropy-code $\bm{c}=Q_{\Lambda}(\bm{y}-\bm{u})$ with entropy models $p_{\bm{c}|\bm{u}}$ and $p_{\bm{c}|\hat{\bm{u}}}$ that are conditioned on their respective shared randomness $\bm{u}$ and $\hat{\bm{u}}$.
@@
In this paper, we investigate how neural compressors may achieve RDP optimality, and what components are necessary for good RDP performance. The RDP coding theorems, while non-constructive, shed light on properties of RDP-optimal compressors. In contrast to the classical rate-distortion function $R(D)=R(D,\infty)$, which is asymptotically achievable by fully deterministic codes, achieving the RDP function may require not only stochastic encoding/decoding but also infinite randomness shared between the encoder and decoder. Devising neural compressors that may achieve RDP optimality at low complexity is an important step in advancing the theory and practice of neural compression. Moreover, infinite shared randomness may not always be available. In settings where shared randomness is limited, or unavailable, we are still interested in schemes that achieve the best possible performance.
@@
Infinite shared randomness has previously proved successful in RDP-oriented compressors such as Theis et al. (2022) that are based on reverse channel coding (RCC) (Theis and Ahmed, 2022; Li and El Gamal, 2018; Li, 2024; Cuff, 2013). RCC enables communication of a sample from a prescribed distribution (e.g., one that is good for RDP) under a limited rate constraint. The performance of RCC is provably near-optimal, but this comes at the cost of high complexity. Moreover, RCC heavily relies on infinite amount of randomness and does not allow for limited or zero randomness. We seek to develop methods with much lower complexity and allow for zero, limited, or infinite amount of randomness.
@@
In the classical rate-distortion framework, recent work has investigated whether neural compressors are optimal, where vector quantization (VQ) (Gersho and Gray, 2012) is known to be optimal, but suffers high complexity. In Lei et al. (2025), it was shown that lattice transform coding (LTC), which uses lattice quantization (LQ) in the latent space, can achieve a performance close to VQ at significantly lower complexity; see also Zhang and Wu (2023); Kudo et al. (2023). VQ and LTC provide near-optimal RD performance due to the efficient way that they pack the space. However, it is not clear at first glance whether VQ-like coding is good for the RDP setting, where randomized reconstructions are often required (Tschannen et al., 2018) to satisfy the perception constraint. It is further unclear how randomness should be incorporated with quantization in a way that is RDP-optimal.
@@
Dithering is a common method of injecting randomness in the quantization process. Classically, a shared random dither can help the quantization noise admit desired statistical properties, and has applications ranging from universal quantization (Ziv, 1985; Zamir et al., 2014) to practical training methods for neural compression (Ballé et al., 2020). For the RDP setting, using dithering to introduce randomness has been investigated recently; Theis and Agustsson (2021) show on a simple source how dithered SQ benefits RDP.
@@
In this paper, we introduce randomness, shared and private, into neural compressors via architectures that build on LTC (Fig. 1), and investigate the roles that randomness and quantization play in how neural compressors perform for the RDP tradeoff while managing complexity. We study LTC where randomness takes the form of a random dither vector. When the randomness is private (i.e., none shared), the dither is only added at the decoder (Fig. 1(b)). When the randomness is shared, the dither is subtracted at the encoder, and added back at the decoder (Figs. 1(c), 1(d)). This work establishes a unified perspective between transform coding, lattice coding, and randomized coding, identifying the roles each play when used together. Our contributions are the following.
@@
    We propose lattice transform coding (LTC) with infinite or no shared randomness, using a shared or private dither respectively. We provide intuition on why shared dithering is beneficial, and describe how such models can be trained in Sec. 3. We then propose a novel dithering scheme that allows for finite randomness to be shared between the encoder and decoder, by using a discrete dither defined via nested lattices. This scheme interpolates between private and shared dithering, enabling control over the rate of shared randomness.
@@
    We theoretically analyze LTC with private and shared dithering on the Gaussian source in Sec. 4. For the latter, we show that the RDP function $R(D,P)$ can be achieved. For the former, under a squared-Wasserstein perception constraint of $P=0$, we show that $R_{p}(D,0)$ is achievable, where $R(D,0)<R_{p}(D,0)<R(\nicefrac{{D}}{{2}},\infty)$. This is strictly worse than the RDP function but strictly better than the “twice-the-distortion” rate-distortion limit.
@@
    We empirically study their performance on synthetic sources with finite-dimensional lattices in Sec. 5. We further show on real-world sources that performance improves with increased shared randomness and better lattice packing efficiency.
@@ 2 Background and Related Work
2.1 Neural Compression for RDP
Most neural compressor designs that account for perception are derivative of the nonlinear transform coding (NTC) setup (Ballé et al., 2020). These models are parameterized by analysis transform $g_{a}$, synthesis transform $g_{s}$, and entropy model $p_{\hat{\bm{y}}}$; see Fig. 1(a). To compress a source $\bm{x}$, the encoder computes the latent $\bm{y}=g_{a}(\bm{x})$, which gets scalar quantized via rounding. The codeword or quantized latent $\hat{\bm{y}}=Q_{\Lambda}(\bm{y})$ is then entropy coded using an entropy model $p_{\hat{\bm{y}}}$. The decoder provides the reconstruction $\hat{\bm{x}}=g_{s}(\hat{\bm{y}})$. The model is trained end-to-end via
@@ 2.1 Neural Compression for RDP
where $\theta$ denotes the parameters of the codec $(g_{a},g_{s},p_{\hat{\bm{y}}})$, $\lambda_{1},\lambda_{2}\geq 0$ are parameters that select a regime of the RDP tradeoff, and the randomness is over $\bm{x}$ and potentially the codec. Many state-of-the-art methods (Tschannen et al., 2018; Agustsson et al., 2019; Mentzer et al., 2020; Muckley et al., 2023; He et al., 2022; Zhang et al., 2021) optimize (2), primarily differing in the choice of distortion function $\Delta$ and the way the statistical divergence $\delta$ is estimated in practice. Typically, $\delta$ is estimated using generative adversarial networks (Goodfellow et al., 2014), which estimate $\delta$ using an adversarial loss involving an auxiliary function parameterized as a discriminator neural network. The use of randomness (shared or not) in these methods has not always been consistent, nor fully explored in its relation to RDP optimality. While Blau and Michaeli (2019) add uniform noise to the quantized latent, Tschannen et al. (2018); Agustsson et al. (2023) concatenates noise to the quantized latent, and Mentzer et al. (2020); Muckley et al. (2023) do not use randomness at all. Zhang et al. (2021) use (shared) dithered scalar quantization with NTC, but do not explore why or how it is good for RDP. In contrast, we study lattice quantization with dithering under the settings of infinite shared, limited shared, and no shared randomness, and show that one needs both the improved packing efficiency of lattices along with shared lattice dithering to achieve best performance.
While there exist a few RDP-oriented neural compressors that do not fit the NTC framework (Theis et al., 2022; Yang et al., 2024; Yang and Mandt, 2024) and instead leverage diffusion models, our work focuses on NTC-style neural compressors for RDP, as they remain the most pervasive type of neural compressor in use, and do not suffer from the higher complexity of RCC or diffusion models.
2.2 Information-Theoretic Analysis of RDP
@@
RDP Theory and Coding Theorems. The RDP function in (1) was originally proposed by Blau and Michaeli (2019); see also Matsumoto (2018); Li et al. (2011); Saldi et al. (2015). (1) is a purely informational quantity, i.e., simply a function of the source $P_{X}$, and thus on its own it does not have a meaning as a fundamental limit of compression without a corresponding coding theorem describing it as such. The first RDP coding theorem was provided by Theis and Wagner (2021), who showed that (1) is achievable via a coding scheme with shared randomness, as well as the converse, i.e., that no compressor can outperform (1). Under perfect realism ($P=0$), Wagner (2022) further characterizes the fundamental RDP limit when only $R_{c}$ bits per sample of shared randomness is allowed, which only coincides with (1) when $R_{c}=\infty$; any less randomness results in a strictly worse fundamental limit. This establishes the necessity of infinite shared randomness to achieve (1). Salehkalaibar et al. (2024); Niu et al. (2023) study a slightly different setting of conditional RDP. Hamdi et al. (2024a) show that randomized encoders do not help in the private randomness setting. Li et al. (2011) use dithered lattice quantization to show achievability of the RDP function for $P=0$; in contrast, we show this to be true for general $P$ and also analyze the private randomness case. Chen et al. (2022) study the RDP tradeoff when the perception constraint is operationally measured in the strong or weak-sense [^2] On vectors $\bm{x},\hat{\bm{x}}\in\mathbb{R}^{n}$, a strong-sense perception constraint denotes $\delta(P_{\bm{x}},P_{\hat{\bm{x}}})<P$, whereas a weak-sense constraint denotes $\delta(P_{\bm{x}_{i}},P_{\hat{\bm{x}}_{i}})<P,\forall i=1,\dots,n$.. Under weak-sense, it was shown that $R(D,P)$ is achievable without shared randomness, whereas under strong-sense, shared randomness is necessary, agreeing with Wagner (2022). In our work, we focus on strong-sense, since that is typically how the perception is measured in practice (i.e., in (2)). In particular, we focus on strong-sense Wasserstein, since that aligns with evaluation metrics used in practice, such as FID.
Coding theorems, while useful for theoretical analysis, are typically not constructive and/or not practical (e.g., due to high complexity). They do, however, provide insights on structures that may be useful or even necessary for optimality, such as (shared) randomness or VQ-like codebooks, that we discuss in the next section. In contrast, we propose a concrete scheme that is both optimal and low complexity.
@@
The benefits of (potentially shared) randomness have also been discussed outside the context of coding theorems. It has been known that randomized decoders are necessary to achieve perfect perceptual quality (Tschannen et al., 2018). Theis and Agustsson (2021) illustrate how quantizers can benefit from shared randomness on a toy circle source. Similarly, Zhou and Tian (2024) demonstrate the benefit of staggered scalar quantizers can use limited shared randomness to improve performance on the circle. In contrast, our work presents a more general approach for infinite, limited, and no shared randomness with lattice quantizers that empirically shows their benefits and is provably optimal on Gaussians.
Reverse Channel Coding. Reverse channel coding (RCC), also known as channel simulation, devises schemes for sending a sample from a prescribed distribution with a rate constraint (Cuff, 2013; Li, 2024), and typically assumes shared randomness; a more comprehensive introduction can be found in Appendix A. Since RDP requires randomized reconstructions, RCC can be applied to RDP, and in fact is a technique to show achievability of RDP under infinite shared randomness (Theis and Wagner, 2021). The particular RCC scheme is that of Li and El Gamal (2018), which samples the same random sequence $\hat{X}_{1},\hat{X}_{2},...\sim P_{\hat{X}}$ at the encoder and decoder (via shared randomness), where $P_{\hat{X}}$ is the optimal reconstruction distribution under (1). To compress $X$, an index $K$ is chosen under a certain criterion, and entropy-coded; the decoder returns $\hat{X}_{K}$; see Def. 27. It was shown in Lei et al. (2022) that this criterion for classical rate-distortion finds the $\hat{X}_{K}$ that is minimum distance to $X$, plus a random regularizer that depends on the rate. In the RDP setting, Prop. 31 reveals a similar minimum-distance-based encoding is performed. Thus, when taking distortion into account, RCC can be seen as a sort of randomized VQ that finds the closest codeword (which are random). Despite its optimality for RDP, RCC is of high complexity (exponential in rate and dimension), and requires infinite shared randomness. What RCC indicates, however, is that a good scheme for RDP should have VQ-like packing efficiency (i.e., good for distortion) combined with random codewords that follow the right distribution (i.e, good for perception). In our work, lattice quantization allows for the VQ-like packing efficiency with low complexity, while the dithering allows the codewords to be stochastic. We note that Sriramu et al. (2024) recently lowered the complexity of RCC schemes via polar codes; however, this approach is limited to discrete data, whereas our work assumes continuous data.
2.3 Lattice and Dithered Quantization
@@
Dithered quantization has been explored in the scalar case (Ballé et al., 2020; Agustsson and Theis, 2020; Zhang et al., 2021) for various reasons including universal quantization (in which the reconstruction is uniformly distributed), but its impact on optimality for the RDP setting has not been fully explored. In contrast, we propose dithering combined with LQ and analyze its effect on RDP performance.

### v2

@@ Abstract
Recent efforts in neural compression have focused on the rate-distortion-perception (RDP) tradeoff, where the perception constraint ensures the source and reconstruction distributions are close in terms of a statistical divergence. Theoretical work on RDP describes properties of RDP-optimal compressors without providing constructive and low complexity solutions. While classical rate distortion theory shows that optimal compressors should efficiently pack space, RDP theory additionally shows that infinite randomness shared between the encoder and decoder may be necessary for RDP optimality. In this paper, we propose neural compressors that are low complexity and benefit from high packing efficiency through lattice coding and shared randomness through shared dithering over the lattice cells. For two important settings, namely infinite shared and zero shared randomness, we analyze the RDP tradeoff achieved by our proposed neural compressors and show optimality in both cases. Experimentally, we investigate the roles that these two components of our design, lattice coding and randomness, play in the performance of neural compressors on synthetic and real-world data. We observe that performance improves with more shared randomness and better lattice packing.
@@ 1 Introduction
$\mathrm{s.t.}$ $\operatorname{\mathbb{E}}_{P_{X}P_{\hat{X}|X}}[\operatorname{\Delta}(X,\hat{X})]\leq D,\quad\delta(P_{X},P_{\hat{X}})\leq P,$
@@
(a) Lattice transform coding (LTC); deterministic.
(b) Private-dither LTC (PD-LTC); $R_{c}=0$.
(c) Shared-dither LTC (SD-LTC); $R_{c}=\infty$.
(d) Quantized-shared-dither LTC (QSD-LTC); finite $R_{c}$.
@@
Figure 1: Lattice transform coding (LTC) with $R_{c}$ bits of (shared) randomness using dithering; $\bm{u}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda))$ and $\bm{u}_{f}\sim\mathrm{Unif}(\mathcal{V}_{0}(\Lambda_{f}))$ are continuous, and $\hat{\bm{u}}\sim\mathrm{Unif}(|\Lambda/\Lambda_{f}|)$ is discrete, where $\Lambda,\Lambda_{f}$ are nested lattices. LTC/PD-LTC entropy-code $Q_{\Lambda}(\bm{y})$ with likelihoods $p_{\hat{\bm{y}}}$. SD-LTC and QSD-LTC entropy-code $Q_{\Lambda}(\bm{y}-\bm{u})$ and $Q_{\Lambda}(\bm{y}-\hat{\bm{u}})$ with likelihoods $p_{\bm{c}|\bm{u}}$ and $p_{\bm{c}|\hat{\bm{u}}}$, respectively.
@@
In this paper, we investigate how neural compressors may achieve RDP optimality, and what components are necessary for good RDP performance. The RDP coding theorems, while non-constructive, shed light on properties of RDP-optimal compressors. In contrast to the classical rate-distortion function $R(D)=R(D,\infty)$, which is asymptotically achievable by fully deterministic codes, achieving the RDP function may require not only stochastic encoding/decoding but also infinite randomness shared between the encoder and decoder. Devising neural compressors that may achieve RDP optimality at low complexity is an important step in advancing the theory and practice of neural compression. Moreover, infinite shared randomness may not always be available. In settings where shared randomness is limited, or unavailable, we are still interested in the best possible schemes.
@@
Infinite shared randomness has previously proved successful in RDP-oriented compressors such as Theis et al. (2022) that are based on reverse channel coding (RCC) (Theis and Ahmed, 2022; Li and El Gamal, 2018; Cuff, 2013). RCC enables communication of a sample from a prescribed distribution (e.g., one that is good for RDP) under a limited rate constraint. The performance of RCC is provably near-optimal, but this comes at the cost of high complexity. Moreover, RCC heavily relies on infinite amount of randomness and does not allow for limited or zero randomness. We seek to develop methods with much lower complexity and allow for zero, limited, or infinite amount of randomness.
@@
In the classical rate-distortion framework, recent work has investigated whether neural compressors are optimal, where vector quantization (VQ) (Gersho and Gray, 2012) is known to be optimal, but suffers high complexity. In Lei et al. (2025), it was shown that lattice transform coding (LTC), which uses lattice quantization (LQ) in the latent space, can achieve a performance close to VQ at significantly lower complexity; see also Zhang and Wu (2023); Kudo et al. (2023). VQ and LTC provide near-optimal RD performance due to high space-packing efficiency. However, it is not clear at first glance whether VQ-like coding is good for the RDP setting, where randomized reconstructions are often required (Tschannen et al., 2018) to satisfy the perception constraint. It is further unclear how randomness should be incorporated with quantization in a way that is RDP-optimal.
@@
Dithering is a common method of randomizing a quantizer. Classically, a shared random dither can help quantization noise admit desired statistical properties, and has applications such as universal quantization (Ziv, 1985), and practical training methods for neural compression (Ballé et al., 2020). For the RDP setting, using dithering to introduce randomness has been investigated recently; Theis and Agustsson (2021) show on a simple source how dithered scalar quantization (SQ) benefits RDP.
@@
In this paper, we introduce randomness, shared and private, into neural compressors via architectures that build on LTC (Fig. 1), and investigate the roles that randomness and quantization play in how neural compressors perform for the RDP tradeoff while managing complexity. We study LTC where randomness takes the form of a random lattice dither vector. When the randomness is private (i.e., none shared), the dither is only added at the decoder (PD-LTC; Fig. 1(b)). When the randomness is shared, the dither is subtracted at the encoder, and added back at the decoder (SD-LTC and QSD-LTC; Figs. 1(c), 1(d) respectively). This work unifies transform coding, lattice coding, and randomized coding, identifying the roles each play when used together for RDP. Our contributions are the following.
@@
    We propose LTC with infinite or no shared randomness, using a shared dither (SD-LTC) or private dither (PD-LTC) respectively, and describe the benefits of the former. We then propose a discrete dithering scheme defined via nested lattices that allows for finite randomness to be shared between the encoder and decoder. This scheme, QSD-LTC, interpolates between PD- and SD-LTC, enabling control over the rate of shared randomness.
@@
    We theoretically analyze and show optimality of SD-LTC and PD-LTC on the Gaussian source with squared 2-Wasserstein perception and asymptotic blocklength in Sec. 4. For the former (where infinite shared randomness is available), we use the sphere-like behavior of lattice cells and the AWGN-like behavior of dithered LQ to show that SD-LTC achieves the RDP function $R(D,P)$. For the latter (where no shared randomness is available), we use lattice Gaussian coding methods to show that under a perception constraint of $P=0$, PD-LTC achieves $R(\nicefrac{{D}}{{2}},\infty)$, which coincides with the fundamental RDP limit under no shared randomness.
@@
    We empirically study PD-LTC, SD-LTC, and QSD-LTC performance on synthetic and real-world sources in Sec. 5. We verify our theory and show that RDP performance improves with increased shared randomness and better lattice packing efficiency.
@@ 2 Background and Related Work
Neural Compression for RDP. Most neural compressor designs that account for perception are derivative of the nonlinear transform coding (NTC) setup (Ballé et al., 2020). These models are parameterized by analysis transform $g_{a}$, synthesis transform $g_{s}$, and entropy model $p_{\hat{\bm{y}}}$; see Fig. 1(a). To compress a source $\bm{x}$, the encoder computes the latent $\bm{y}=g_{a}(\bm{x})$, which gets scalar quantized via rounding. The codeword or quantized latent $\hat{\bm{y}}=Q_{\Lambda}(\bm{y})$ is then entropy coded using an entropy model $p_{\hat{\bm{y}}}$. The decoder provides the reconstruction $\hat{\bm{x}}=g_{s}(\hat{\bm{y}})$. The model is trained end-to-end via
@@
where $\theta$ denotes the parameters of the codec $(g_{a},g_{s},p_{\hat{\bm{y}}})$, and $\lambda_{1},\lambda_{2}\geq 0$ control the RDP tradeoff. Many state-of-the-art methods (Tschannen et al., 2018; Agustsson et al., 2019; Mentzer et al., 2020; Muckley et al., 2023; He et al., 2022; Zhang et al., 2021) optimize (2), primarily differing in the choice of $\Delta$ and the way $\delta$ is estimated in practice, which is typically done using an adversarial loss involving a discriminator neural network. The use of randomness (shared or not) in these methods has not always been consistent, nor fully explored in its relation to RDP optimality. While Blau and Michaeli (2019) add uniform noise to the quantized latent, Tschannen et al. (2018); Agustsson et al. (2023) concatenate noise to the quantized latent, and Mentzer et al. (2020); Muckley et al. (2023) do not use randomness at all. Zhang et al. (2021) use (shared) dithered SQ with NTC, but do not explore why or how it is good for RDP. In contrast, we study LQ with dithering under varying levels of shared randomness, and show that one needs both the improved packing efficiency of lattices along with shared lattice dithering to achieve best performance. While there exist a few RDP-oriented neural compressors that do not fit the NTC framework (Theis et al., 2022; Yang et al., 2024; Yang and Mandt, 2024) and instead leverage diffusion models, our work focuses on NTC-style neural compressors for RDP, as they remain the most pervasive type of neural compressor in use, and do not suffer from the higher complexity of RCC or diffusion models.
@@
Information-Theoretic Analysis of RDP. The RDP function in (1) was formally introduced and widely adopted following Blau and Michaeli (2019); see also earlier related work by Matsumoto (2018); Li et al. (2011); Saldi et al. (2015). (1) is a purely informational quantity, i.e., a function of the source $P_{X}$, and thus does not have a meaning as a fundamental limit of compression without a corresponding coding theorem describing it as such. The first RDP coding theorem was provided by Saldi et al. (2015), who show that for perfect perception ($P=0$), (1) is achievable under infinite shared randomness, and conversely that no compressor can outperform (1). For general $P$, Theis and Wagner (2021) establish optimality (i.e., achievability and converse) of (1) when infinite shared randomness is available. Saldi et al. (2015) further characterizes the fundamental RDP limit at $P=0$ when only $R_{c}$ bits per sample of shared randomness is allowed, which only coincides with (1) when $R_{c}=\infty$; a smaller $R_{c}$ results in a strictly worse fundamental limit (see also Wagner (2022)). This establishes the necessity of infinite shared randomness to achieve (1). Li et al. (2011) use dithered LQ to show achievability of (1) for $P=0$; in contrast, we show this to be true for general $P$ and also analyze the $R_{c}=0$ case, which requires a new set of proof techniques based on lattice Gaussian coding (Ling and Belfiore, 2014). Under private randomness ($R_{c}=0$), Yan et al. (2021) establish $R(\nicefrac{{D}}{{2}},\infty)$ as the fundamental limit for $P=0$; Hamdi et al. (2024a) show that randomized encoders do not help. Chen et al. (2022) study the RDP tradeoff when the perception constraint is strong- or weak-sense [^2] On vectors $\bm{x},\hat{\bm{x}}\in\mathbb{R}^{n}$, strong-sense denotes $\delta(P_{\bm{x}},P_{\hat{\bm{x}}})<P$; weak-sense denotes $\delta(P_{\bm{x}_{i}},P_{\hat{\bm{x}}_{i}})<P,\forall i$.. Under weak-sense, it was shown that $R(D,P)$ is achievable without shared randomness, whereas under strong-sense, shared randomness is necessary, agreeing with Saldi et al. (2015); Wagner (2022). In our work, we focus on strong-sense, since that is typically how the perception is measured in practice (i.e., in (2)). In particular, we focus on strong-sense Wasserstein, since that aligns with $\delta$’s and evaluation metrics chosen in practice such as Fréchet inception distance.
Coding theorems use schemes that are typically not constructive (e.g. random coding) and/or not practical (e.g., high complexity RCC schemes). They do, however, provide insights on structures that may be useful or even necessary for optimality. In addition to the necessity of infinite shared randomness to achieve (1), we show in Sec. A how the RCC scheme of Theis and Wagner (2021) implies that a good RDP compressor should behave like a randomized VQ: it should have VQ-like packing efficiency (i.e., good for distortion) combined with random codewords that follow the right distribution (i.e, good for perception). In our work, the former is handled with LQ (at low complexity), while the latter is handled with dithering; our schemes are further shown optimal on Gaussians.
@@
The benefits of (potentially shared) randomness have also been discussed outside the context of coding theorems. Tschannen et al. (2018) show how randomized decoders are necessary to achieve perfect perceptual quality. Theis and Agustsson (2021) illustrate how quantizers benefit from shared randomness on a toy circle source. Similarly, Zhou and Tian (2024) demonstrate how staggered SQ can use limited shared randomness to improve performance on the circle. In contrast, our work presents a more general approach for infinite, limited, and no shared randomness with LQ that empirically shows its benefits and is provably optimal on Gaussians.